OpenStudy (stormswan):

Please help me, i will give a medal and fan. Bruce added some chlorine to the water in a pool. The chlorine evaporated at a fixed rate every week. The table below shows the amount of chlorine f(n), in ounces, that was left in the pool after n weeks: n f(n) 1 54 2 27 3 13.5 4 6.75 Which function best shows the relationship between n and f(n)?

OpenStudy (stormswan):

f(n) = 108(0.5)^n-1 f(n) = 54(0.5)^n f(n) = 108(0.5)^n+1 f(n) = 54(0.5)^n-1

OpenStudy (ikram002p):

what do u think /.?

OpenStudy (stormswan):

i honestly don't even know. thats why i asked for help

OpenStudy (ikram002p):

since its fixed rate , find rate of first two terms rate=27/54=0.5

OpenStudy (ikram002p):

so f(n)=first term *( rate )^n-1

OpenStudy (stormswan):

(0.5) right?

OpenStudy (ikram002p):

?

OpenStudy (stormswan):

idk what I'm doing

OpenStudy (ikram002p):

use the formula of f(n)=first term *( rate )^n-1 first term =54 rate =0.5 thn f(n) = 54(0.5)^n-1

OpenStudy (stormswan):

and thats it?

OpenStudy (ikram002p):

yeah !

OpenStudy (stormswan):

oh! ok. thank you!

OpenStudy (ikram002p):

np :)

OpenStudy (anonymous):

so what was the answer @Stormswan?

OpenStudy (stormswan):

@xfordboy97x f(n) = 54(0.5)^n-1